OpenStudy (pphalke):

Randomly draw a card from a deck of 52 cards. Find the probability of drawing a diamond card OR a face card.

OpenStudy (mathstudent55):

You need to figure out how many cards are diamond or face. This is a way of doing it: Find out the number of diamond cards Find out the number of face cards. Add the two numbers above together. This is not your final number. The cards that are both diamond and face were counted twice, so this number is larger than it should be. We need to fix that. Now find the number of cards that are both diamond and face. Subtract this number from the sum above. This adjusts you number above to have each card only counted once. This is now the number of cards that are face or diamond. Divide this final number by the number of cards in a deck.

OpenStudy (pphalke):

This is a bit confusing

OpenStudy (sql_server):

Let me give a try

OpenStudy (sql_server):

Tell me first how many diamond cards are there in a deck?

OpenStudy (dumbcow):

12 face cards (4 jack,4queen,4 king) 10 other diamond cards (A,2,3,4,5,6,7,8,9,10) 22 total prob = 22/52 = 11/26